the number of dogs licensed in a particular city this year is 1300. this is a 5% increase over last year.how many dog licenses were on record last year? round to the nearest whole number.
Answer by ewatrrr(24785)


Hi
Let x represent many dog licenses were on record last year
Question states***
(1 + .05)x = 1300
Solving for x
x = 1300/1.05
x = 1238, the dog licenses on record last year
CHECKING our Answer***
1238 + .05*1238 = 1238 + 62 = 1300